OpenStudy (catlover12345678):

solve for x 3x+1/14=2 x/8 please help

RhondaSommer (rhondasommer):

is it this? \[x~\frac{3x+1}{14}=2\frac{x}{8}\]

OpenStudy (catlover12345678):

is it -2/77